OpenStudy (daniel14):

what number would you subtract from each side of the inequality to solve to solve: y+3/8(greater than sign here)16 A. -16 B. -3/8 C. 3/8 D. 16

OpenStudy (daniel14):

I am so lost

OpenStudy (jdoe0001):

ok hmmm say if you had x + 3 = 25 what would "x" equals to?

OpenStudy (daniel14):

22

OpenStudy (jdoe0001):

ahah! 22 well... how did you get 22 though?

OpenStudy (daniel14):

are u really asking lol haha umm maybe something called easy math hey? hehe

OpenStudy (daniel14):

whats the next step?

OpenStudy (daniel14):

@jdoe0001

OpenStudy (jdoe0001):

well seems all you did was "subtract" the -3 from both sides x+3-3 = 25 -3 x +0 = 22 x = 22 well.. you'd do the same thing with the equation above \(\large { y+\cfrac{3}{8}> 16\implies y\cancel{+\cfrac{3}{8}-\cfrac{3}{8}}> 16-\cfrac{3}{8}\implies y> 16-\cfrac{3}{8} }\)